Container A measuring 2m by 70cm by 1.5m is 4/5 filled with water.The water in container A is poured into container B with a base of 150cm by 175cm.What is the height of the water in container B in metres?
1
Expert's answer
2011-10-07T11:27:56-0400
The volume of water in container A is: Va = (4/5)*(2x0.7x1.5) = 1.68 [kL]
The volume of water in container B is: Vb = 1.5x1.75xH = 2.625H, where H is unknown height.
It must be the same as Va: va = Vb. So, 2.625H = 1.68 and H = 1.68/2.625 = 0.64m.
So, the height of the water in container B is 0.64 metres.